The ages of Paul and Judy total 27 years
Judy's age plus twice Paul' age is 43
First translate the words into an algebra equation

Answer by Alan3354(69443)

The ages of Paul and Judy total 27 years
Judy's age plus twice Paul' age is 43
First translate the words into an algebra equation
----------------
Use P for Paul's age and J for Judy's age. It saves typing.
The 1st sentence says P + J = 27. So far so good.
Then the 2nd sentence:
J + 2*P = 43 See that?
Now we have 2 eqns in 2 unknowns, so we can solve for P and J (not PB & J, tho)
---------
J + P = 27 EQN 1
J + 2P = 43 EQN 2
Subtract EQN 1 from EQN 2
0 + P = 16
So Paul is 16
27 - 16 = 11, so
Jane is 11.
That's it.